how do you solve the quadratic equation 2x^2=-(7x+3)

Answer by checkley77(12844)

2x^2=-(7x+3)
2x^2+7x+3=0
(2x+1)(x+3)=0
2x+=1=0
2x=-1
x=-1/2 ans.
x+3=0
x=-3 ans.